OpenStudy (blockcolder):

Prove that \(\sqrt{n+1}-\sqrt{n}=O(\sqrt{n}\). No idea how to start. Hints please?

OpenStudy (anonymous):

Multiply by a creative use of "1." Hint, think of the conjugate of \[\sqrt{n+1} - \sqrt{n}\]

OpenStudy (anonymous):

Also, write out the definition of Big O of a function again. Write it here so everyone can see and be on the same page, sometimes books use slightly different definitions

OpenStudy (anonymous):

@blockcolder What is the "O" on the right hand side of the equation?

OpenStudy (blockcolder):

f(x) is O(g(x)) if \(|f(x)| \leq C|g(x)|\) for sufficiently large x.

OpenStudy (blockcolder):

Going by @domu 's advice, I got \(\large \sqrt{n+1}-\sqrt{n}=\frac{1}{\sqrt{n+1}-\sqrt{n}}\) I suppose the next step is to bound that last one above?

OpenStudy (anonymous):

Said otherwise \[\lim_{x \rightarrow \infty} \frac{ f(x) }{ g(x) } =C\]

OpenStudy (anonymous):

Yup yup, your on the right path. Though that " - " sign in the denominator should be a " + " Everything else is looking good so far!

OpenStudy (blockcolder):

Does this look ok? \[\frac{1}{\sqrt{n+1}+\sqrt{n}}<\frac{1}{2\sqrt{n}}<\frac{1}{\sqrt{n}}<\sqrt{n}\] since we're considering large n.

OpenStudy (anonymous):

Looks good to me! Another way would be to note that the left hand side tends to 0 as n gets large. Your answer is much more concrete though and I like it : ) Nice work

OpenStudy (blockcolder):

Thanks for the help.